In the Word Pencils, how many combos can you make if the "N" has to be the first letter and P and I have to be next to each other.

Answer by stanbon(75887)

In the Word Pencils, how many arrangements can you make if the "N" has to be the first letter and P and I have to be next to each other.
-----
P and I act as one letter with 2 arrangements.
So there are 6 letters and one of the letters has 2 arrangements.
----
1st letter is "n".
------
# of arrangements: 1*2*5! = 2*120 = 240
=====================
Cheers,
Stan H.
---